# Realistic Cost to Execute Practical Quantum Circuits using Direct Clifford+T Lattice Surgery Compilation

### v2

@@ 2.3.3. Routing and Parallelism using the Wave Pipeline
Our modified version of LSC detects parallel instructions from the logical circuit and attempts to schedule as many instructions on each slice as possible while avoiding contention. The new wave pipeline, made accessible through the command line option -P wave, uses a greedy scheduling algorithm within the second stage of compilation to route and slice logically parallel operations in “waves” according to both their logical dependency graph and constraints imposed by the layout. The wave pipeline constructs a directed acyclic graph (DAG) for the entire layout-independent computation, and includes an As-Soon-As-Possible scheduler that takes routing conflicts into account. (See Algorithm 1 for details.) One downside of this approach is the abandonment of the streaming philosophy underlying the original LSC. While we imagine that a windowed-DAG methodology could be used in order to obtain parallelism while not sacrificing the ability to stream instructions, we reserve this for later work.
@@ 2.3.4. Local Compilation Layer
TwoPatchMeasure
Measures the joint XX/ZZ operators of two vertically/horizontally-adjacent initialized tiles
2 1
BellPrepare
Initializes a Bell state on two adjacent uninitialized tiles
2 1
BellMeasure
Performs a destructive Bell basis measurement on two adjacent (initialized) tiles and makes uninitialized
2 1
ExtendSplit
A patch extension followed by a split operation
2 1
MergeContract
A merge operation followed by a patch contraction
2 1
Move
A patch extension followed by a patch contraction
2 1
Patch Contraction
Contracts one initialized two-tile patch into a one-tile patch while preserving the encoded state
2/1 0
Patch Extension
Extends one initialized one-tile patch into a two-tile patch while preserving the encoded state
1/2 1
@@ 2.3.5. Outputs of the Lattice Surgery Compiler
In addition to the outputs available in the original LSC, we have implemented the --printlli sliced option which yields a line-by-line (slice-by-slice) list of comma-separated LLI with their associated patch IDs. In the case of the BellBasedCNOT LLI, the further compilation into local instructions over multiple slices appears in brackets along with the pair of tiles on which each local instruction acts. We have additionally enabled the LSC to compute and output certain useful statistics such as the active volume of the logical computation using the slice visitation methodology described in Ref. (Watkins et al., 2023).
@@ 3. Performance of the Lattice Surgery Compiler
Now that we have described our compilation methodology and its implementation in the LSC, we demonstrate its performance on random circuits. The performance metrics we consider are compilation time and ability to preserve logical parallelism in the fault-tolerant layer. We also provide insight into the space-time trade-offs that exist within the EDPC layout family. Figures that are plotted against the number of logical qubits have a constant circuit depth and vice versa. To fix circuit depth, we track the number of gates with support on each qubit and add random gates until any qubit has achieved the desired depth. Our circuits contain random CNOT gates, random T gates, or a 50/50 mixture of these. We note that, since our methodology heavily relies on gate teleportation circuits, real circuits boil down to CNOT and Hadamard gates. As such, we focus on CNOT gates to demonstrate trade-offs related to routing and parallelism, and mix in T gates to capture the impact of the requirement to route to boundary tiles to fulfill MagicStateRequests and YStateRequests as well as the impact of Hadamards (from the corrective S circuit).
@@
In Fig. 5(a), we plot the number of slices in the compiled output against the number of logical qubits $x$. With circuit depth fixed, we find that this quantity scales roughly as $O(\sqrt{x})$. This is similar to the worst-case expectation from Ref. (Beverland et al., 2022a) (mentioned previously in Sec. 2.2). Unsurprisingly, the number of slices increases with the proportion of T gates. We suspect that the greater variance in circuits containing T gates results from jumps in the number of tiles reserved for magic states as the layout increases in size. In Fig. 5(b), we plot the compilation time per gate against the number of logical qubits $x$, with circuit depth fixed. The high levels of contention in our randomly generated circuits have a substantial impact on the compile term performance of our implementation. The dominant cost of our implementation comes from the A* routing algorithm, which scales as $O(x\log x)$ in the worst case. This worst case is realized when no paths can be found by A*, and this occurs precisely when there are routing conflicts. To add insult to injury, if a gate cannot be scheduled in slice $t$, the router is rerun for that gate in slice $t+1$. Consequently, high levels of contention cause the A* router to be called $O(\sqrt{x})$ times for each gate, according to the results in Fig. 5(a). Combining these observations, we expect our compile time per gate to scale as $O(x^{1.5}\log x)$ in high contention scenarios. This prediction is supported by Fig. 5(b), since we see approximately quadratic scaling for circuits with up to 250 logical qubits. We also see that the leading coefficients of our best fit power laws increase with the proportion of T gates due to those gates being compiled into several gates according to our strategy explained in Sec. 2.2.
@@
Figure 5. Scaling of (a) the number of slices in the compiled output and (b) the compilation time per gate against the number of logical qubits present in randomly generated circuits of CNOT gates, T gates, and a 50/50 mixture of these. The circuit depth is kept fixed at 10. Trend-lines are power-law fits starting from 20 logical qubits.
@@
Figure 6. The number of slices in the compiled output (a) and the total logical space-time volume (b) against the circuit depth in randomly generated circuits of a 50/50 mixture of CNOT and T gates. The number of logical qubits is kept fixed at 32.
@@
The layouts generated using different combinations of the num_lanes and condensed parameters in the EDPC layout family have different ratios of data:ancilla tiles. In Fig. 4(a), (b), and (c), for example, we see that there is respectively a 3:1, 8:1, and 5:4 data:ancilla tile ratio in the bulk. Additionally, where condensed layouts are concerned each logical qubit tile has only two free edges (for 32 logical qubits, which is divisible by 4), otherwise there are four free edges. In Fig 6(a), it is evident that having more routing lanes leads to a lower number of slices, though the benefit appears to be marginal. Also, the condensed option causes a jump in the number of slices in all cases, especially for 1-lane condensed, which yields the highest number of slices by far. In Fig 6(b), we examine the effect of layout on the total logical space-time volume of the circuit [^15] More precisely, we record the total volume reported by the Lattice Surgery Compiler while excluding magic state distillation and storage costs; this value is the volume of the logical computation given a constant rate of magic state replenishment at reserved cells. The space-time cost of magic state distillation and storage will be accounted for in later sections.. One can see that the total space-time volume for 1-lane condensed is virtually the same as that of 1-lane, and that the condensed layouts generally perform better than their un-condensed counter-parts. The conclusion to be drawn is that 1-lane [Fig. 4(a)] and 1-lane condensed [Fig. 4(c)] are comparably good layouts, and that the choice between them will depend on whether one would prefer to save time or space in their computation. We also conclude that the benefit of opening up additional lanes for routing is probably not worth the space cost. Given these results, we have opted to use the 1-lane EDPC layout in our resource analyses in Sec. 5.
@@ 3.1. Outlook for the Lattice Surgery Compiler
In practice, circuits with dozens of qubits and millions of gates can be compiled in hours (see Table 2 in Sec. 5). Larger circuits, especially those involving many qubits, require decomposition into smaller sub-circuits. We found that the performance of the LSC is more than enough to include key sub-circuits of the ground state estimation circuit compilation methodology that we consider in Sec. 5. Future work should allow LSC to link sub-circuits together; we expect that this strategy would allow much larger circuits to be compiled without issue.
@@ 4.1. Usage of the Lattice Surgery Compiler for Resource Estimation
As seen in Fig. 4, the layout specification generated by the LSC using the -L edpc option contains tiles that are reserved for magic states which are replenished at a rate given by the disttime option. As alluded to in Sec. 2, our resource estimation methodology does not assume a 1:1 correspondence between magic state factories and the tiles reserved for magic states in the layout. Instead, we calculate the necessary number of magic state factories using the rate at which the RequestMagicState LLI is found in the sliced output. The distillation cycle time also depends on the specifics of the distillation factory used, which is also calculated downstream of the LSC output and does not depend on the disttime option. We note that parallelism within the fault-tolerant layer (and therefore algorithm run-time) can improve with a higher rate of replenishing magic states, but in turn increases the number of factories needed to keep up with supply (and therefore the spatial footprint of distillation). Conversely, a longer algorithm run-time or a resulting surplus of reserved magic states can increase the minimum required code distance to remain below a logical error rate threshold, increasing the spatial footprint as well. Due to these non-trivial space-time trade-offs, we expect that disttime is a parameter worthy of optimization. In this work, we make the simplifying assumption that magic states are available as quickly as possible given constraints set by the layout by keeping --disttime 1  [^17] Preliminary tests have demonstrated that this parameter does not have a major impact on the final number of slices as long as it remains below a certain cutoff related to the maximum rate at which magic states can be consumed..
@@ 4.2. Cost of Magic State Distillation and Storage
We primarily follow the magic state factory design framework proposed in Ref. (Beverland et al., 2022b). In this framework, the total factory $\mathcal{D}$ is composed of $R$ distillation rounds $r\in\{1,2,\dots,R\}$, each using $C_{r}$ copies of a distillation unit $M_{r}$ operating at code distance $d_{r}$. We assume a negligible cost of state injection and of transferring the output magic states from one level of distillation to the correct positions and code distances to use in the next level. Though each of the mentioned parameters can be optimized, we make multiple a priori decisions regarding them. We assume $R=2$ levels of distillation where $C_{2}=1$, and we assume that $M$ is a 15:1 distillation unit that has been compiled to the same local instruction set that we use in the LSC (see Fig. 9 of Ref (Beverland et al., 2022b)). We constrain $d_{2}$ to be same code distance as that used in the logical part of the computation (i.e. the full code distance). For a given choice of $d_{1}$, we require that $C_{1}$ be the smallest number that yields $P_{1}^{acc}(\geq 15)\geq P_{th}$, where $P_{1}^{acc}(\geq 15)$ is the probability of at least 15 output states being accepted at the first level and $P_{th}$ is a pre-decided threshold (taken to be 0.985 here). We note that some combinations of parameters are immediately ruled out due to $P_{2}^{acc}<P_{th}$, since we require $C_{2}=1$. The final run-time of the distillation factory is $\tau(\mathcal{D})\equiv\lceil\tau_{1}+\tau_{2}\rceil$, where $\tau_{r}$ is the number of full-distance slices required for distillation round $r$ [^18] Where $d_{1}\neq d_{2}$, $\tau_{1}=\frac{d_{1}}{d_{2}}\tau_{2}$ may not be a whole number.. The total spatial cost of the distillation factory is $n(\mathcal{D})\equiv\lceil\max_{r}\left[C_{r}n_{r}\right]\rceil=\lceil\max{\left(C_{1}n_{1},n_{2}\right)}\rceil$, where $n_{r}$ is the number of full-distance tiles required for $M$ at distillation round $r$ [^19] Similarly, we calculate $n_{1}$ as $n_{1}=\frac{2d_{1}^{2}-1}{2d_{2}^{2}-1}n_{2}$. The active volume of the distillation factory is approximated to be $V(\mathcal{D})\simeq C_{1}n_{1}\tau_{1}+n_{2}\tau_{2}$. In our resource analysis script, we treat the factory selection process as a function that receives $d_{1}$ and $d_{2}$ as parameters and outputs $n(\mathcal{D})$, $\tau(\mathcal{D})$, $V(\mathcal{D})$, and the distilled state error rate $P_{T}(\mathcal{D})$. This function is used as a subroutine within the resource optimization process to be described.
@@
The object is to calculate $N$, which is the number of magic states that need to be produced in each distillation cycle to keep up with the consumption rate. We note that, since our distillation factories output one state each distillation cycle, $N$ is equal to the number of factories. To cover the requests made during the first cycle, we require $w=\lceil m(1)/N\rceil$ warm-up distillation cycles. Then, we minimize $N$ with respect to the condition $m(k)\leq N(w+k-1)$ because the number of consumed states after $k$ distillation cycles has to be less than or equal to the number produced as of the $(w+k-1)$th cycle at every $k$. Using the expression for $w$, we relax this condition to $m(k)\leq m(1)+Nk$ and solve for N:
(2) $N\geq\frac{m(k)-m(1)}{k}$
@@
(3) $N=\left\lceil\max_{k}\frac{m(k)-m(1)}{k}\right\rceil.$

### v4

@@ 2.3.3. Routing and Parallelism using the Wave Pipeline
Our modified version of LSC detects parallel instructions from the logical circuit and attempts to schedule as many instructions on each slice as possible while avoiding contention. The new wave pipeline, made accessible through the command line option -P wave, uses a greedy scheduling algorithm within the second stage of compilation to route and slice logically parallel operations in “waves” according to both their logical dependency graph and constraints imposed by the layout. The wave pipeline constructs a directed acyclic graph (DAG) for the entire layout-independent computation, and includes an As-Soon-As-Possible scheduler that takes routing conflicts into account. (See Algorithm 1 for details.) This scheduling algorithm was chosen for simplicity; future versions of the compiler could potentially use alternative scheduling algorithms at this stage. One downside of this approach is the abandonment of the streaming philosophy underlying the original LSC. While we imagine that a windowed-DAG methodology could be used in order to obtain parallelism while not sacrificing the ability to stream instructions, we reserve this for later work.
@@ 2.3.4. Local Compilation Layer
TwoPatchMeasure Measures the joint XX/ZZ operators of two vertically/horizontally-adjacent initialized tiles 2 1
BellPrepare Initializes a Bell state on two adjacent uninitialized tiles 2 1
BellMeasure Performs a destructive Bell basis measurement on two adjacent (initialized) tiles and makes uninitialized 2 1
ExtendSplit A patch extension followed by a split operation 2 1
MergeContract A merge operation followed by a patch contraction 2 1
Move A patch extension followed by a patch contraction 2 1
Patch Contraction Contracts one initialized two-tile patch into a one-tile patch while preserving the encoded state 2/1 0
Patch Extension Extends one initialized one-tile patch into a two-tile patch while preserving the encoded state 1/2 1
@@ 2.3.5. Outputs of the Lattice Surgery Compiler
In addition to the outputs available in the original LSC, we have implemented the --printlli sliced option which yields a line-by-line (slice-by-slice) list of comma-separated LLI with their associated patch IDs. In the case of the BellBasedCNOT LLI, the further compilation into local instructions over multiple slices appears in brackets along with the pair of tiles on which each local instruction acts. We have additionally enabled the LSC to compute and output certain useful statistics such as the active volume of the logical computation using the slice visitation methodology described in Ref. (Watkins et al., 2024).
@@ 3. Performance of the Lattice Surgery Compiler
Now that we have described our compilation methodology and its implementation within the LSC, we demonstrate the compiler’s performance on random circuits. To this end, we consider both the average compilation time per gate and the preservation of logical parallelism within the fault-tolerant layer to be key performance metrics. In addition to these metrics, we provide some insight into the space-time trade-offs that exist within the EDPC layout family. To generate random circuits, we specify a number of logical qubits and then add random gates until a cut-off on the circuit depth has been met. Our random circuits contain either CNOT gates or a 50/50 mixture of CNOT gates and T gates. We note that, since our compilation methodology heavily relies on gate teleportation circuits (see Sec. 2.2.1), real circuits boil down to CNOT and Hadamard gates. As such, we focus on CNOT gates to demonstrate trade-offs related to routing and parallelism, and mix in T gates to capture both the impact of the requirement to route to boundary tiles to fulfill MagicStateRequests and YStateRequests and the impact of Hadamard gates (from the corrective S circuit) in blocking paths. We consider scaling in two different regimes: in one, we hold the circuit depth constant and generate circuits with numbers of logical qubits up to 10,000, and in the other, we hold the number of logical qubits constant and vary the circuit depth up to 500. At each grid point we generate ten circuit samples and compile them using the LSC. Each compilation task was performed on a single core of the Compute and Data Environment for Science (CADES) high-performance computing cluster at Oak Ridge National Laboratory.
@@
In Fig. 5(a), we plot the average number of slices in the compiled output against the number of logical qubits $x$ with circuit depth fixed to 10. We find that this quantity scales a little faster than $O(\sqrt{x})$, which is consistent with the worst-case expectation from Ref. (Beverland et al., 2022a) (mentioned previously in Sec. 2.2) [^15] While a more direct comparison requires uniform gate densities, we find that the average gate density decays weakly ($\sim x^{-0.1}$) with the number of logical qubits in our random circuits.. Unsurprisingly, the number of slices increases with the proportion of T gates, which is reflected in the constant factor. In Fig. 5(b), we plot the compilation time per gate against the number of logical qubits $x$ with circuit depth fixed to 10. The high levels of contention in our randomly generated circuits have a substantial impact on the compile time performance of our implementation. The dominant cost of our implementation comes from the A* routing algorithm, which scales as $O(x\log x)$ in the worst case. This worst case is realized when no paths can be found by A*, and this occurs precisely when there are routing conflicts. To add insult to injury, if a gate cannot be scheduled in slice $t$, the router is re-run for that gate in slice $t+1$. Consequently, according to the results in Fig. 5(a), high levels of contention cause the A* router to be called $O(\sqrt{x})$ times for each gate. Combining these observations, we expect our compile time per gate to scale as $O(x^{1.5}\log x)$ in high contention scenarios. This prediction is supported by Fig. 5(b), since we see a little faster than $O(x^{1.5})$ scaling for CNOT circuits up to 10,000 qubits. In the mixed circuits, the slicing of certain instructions may be limited by the availability of resource states rather than by contention for routes, which may explain the lower exponent in this case. We also see that the leading coefficients of our best fit power laws increase with the proportion of T gates due to those gates being compiled into several gates according to our strategy explained in Sec. 2.2. Interestingly, we find that results from the ground state estimation circuits studied in Sec. 5 overlap with the mixed circuit results in Fig. 5(b).
@@
Figure 5. Scaling of (a) the number of slices in the compiled output and (b) the average compilation time per gate against the number of logical qubits present in randomly generated circuits of CNOT gates and a 50/50 mixture of CNOT gates and T gates. The circuit depth is kept fixed at 10 and data points are averaged over 10 circuit samples. Trend-lines are power-law fits starting from 1000 logical qubits. In (b), results for the practical GSE circuits from Table 3 are included for comparison.
@@
Figure 6. The number of slices in the compiled output (a) and the total logical volume (in tile-slices) (b) against the circuit depth in randomly generated circuits of a 50/50 mixture of CNOT and T gates for a variety of layouts within the EDPC layout family. The number of logical qubits is kept fixed at 100 and data points are averaged over 10 circuit samples.
Figure 7. The total logical active volume (in tile-slices) (a) and the same expressed as a percentage of the total logical volume (b) against the circuit depth in randomly generated circuits of a 50/50 mixture of CNOT and T gates for a variety of layouts within the EDPC layout family. The number of logical qubits is kept fixed at 100 and data points are averaged over 10 circuit samples.
@@
Next, Figs. 6 and 7 provide insight into space-time trade-offs that exist within the EDPC layout family. We expect the slice performance of these layouts to differ because the layouts generated using different combinations of the num_lanes and condensed parameters have different ratios of data:ancilla tiles. In Fig. 4(a), (b), and (c), for example, we see that there is respectively a 3:1, 8:1, and 5:4 data:ancilla tile ratio in the bulk. Additionally, in condensed layouts, logical qubit tiles have only two free edges (as long as the number of logical qubits is divisible by 4); otherwise there are four free edges. In Fig 6(a), it is evident that having more routing lanes leads to a lower number of slices, though the benefit appears to be limited. Also, the condensed option causes a jump in the number of slices in all cases, especially for 1-lane condensed, which yields the highest number of slices by far. In Fig 6(b), we examine the effect of layout on the total logical volume of the circuit [^16] More precisely, we record the total volume (in tile-slices) reported by the Lattice Surgery Compiler while excluding magic state distillation and storage costs; this value is the volume of the logical computation given a constant rate of magic state replenishment at reserved tiles, and is independent of code distance. The real space-time cost of algorithms, including contributions from magic state distillation and storage, will be accounted for in later sections.. One can see that, in this respect, condensed layouts generally perform better than their un-condensed counter-parts. That said, while 1-lane condensed [Fig. 4(c)] is the best-performing layout by this metric, its performance is similar to that of 1-lane [Fig. 4(a)]. To better distinguish between them, we further evaluate the effect of layout parameters on the logical active volume (see Sec. 1 for the definition of this quantity), which has implications on total logical error and, accordingly, code distance and space-time volume (see Sec. 4). In Fig. 7, we plot the logical active volume both as an absolute quantity [Fig. 7(a)] and as a percentage of the total volume [Fig. 7b]. We find that, while 1-lane condensed makes most efficient use of its tile-slices by having the highest active volume percentage, the total active volume for 1-lane is lower, yet still similar to that of 1-lane condensed. Ultimately, we conclude that 1-lane and 1-lane condensed are comparably good layouts, and that the choice between them will depend on whether one would prefer to save time or space in their computation. We also conclude that the benefit of opening up additional lanes for routing is probably not worth the space cost. Given these conclusions, we have opted to use the 1-lane EDPC layout in our resource analyses in Sec. 5.
@@ 3.1. Outlook for the Lattice Surgery Compiler
In practice, circuits with dozens of qubits and millions of gates can be compiled in hours (see Table 3 in Sec. 5). Thus, larger circuits, especially those involving many qubits, may require decomposition into smaller sub-circuits that are compiled separately. While in this article we focus on whole-circuit compilation, we have found that the performance of the LSC is more than enough to compile key sub-circuits of the ground state estimation circuits that we consider in Sec. 5 according to the methodology from Ref. (Kornell and Selinger, 2023). Future work should allow LSC to link these sub-circuits together; we expect that this strategy would allow much larger circuits to be compiled without issue.
@@ 4.1. Usage of the Lattice Surgery Compiler for Resource Estimation
As seen in Fig. 4, the layout specification generated by the LSC using the -L edpc option contains tiles that are reserved for magic states which are replenished at a rate given by the disttime option. As alluded to in Sec. 2, our resource estimation methodology does not assume a 1:1 correspondence between magic state factories and the tiles reserved for magic states in the layout. Instead, we calculate the necessary number of magic state factories using the rate at which the RequestMagicState LLI is found in the sliced output. The distillation cycle time also depends on the specifics of the distillation factory used, which is also calculated downstream of the LSC output and does not depend on the disttime option. We note that parallelism within the fault-tolerant layer (and therefore algorithm run-time) can improve with a higher rate of replenishing magic states, but in turn increases the number of factories needed to keep up with supply (and therefore the spatial footprint of distillation). Conversely, a longer algorithm run-time or a resulting surplus of reserved magic states can increase the minimum required code distance to remain below a logical error rate threshold, increasing the spatial footprint as well. Due to these non-trivial space-time trade-offs, we expect that disttime is a parameter worthy of optimization. In this work, we make the simplifying assumption that magic states are available as quickly as possible given constraints set by the layout by keeping --disttime 1181818Preliminary tests have demonstrated that this parameter does not have a major impact on the final number of slices as long as it remains below a certain cutoff related to the maximum rate at which magic states can be consumed..
@@ 4.2. Cost of Magic State Distillation and Storage
We primarily follow the magic state factory design framework proposed in Ref. (Beverland et al., 2022b). In this framework, the total factory $\mathcal{D}$ is composed of $R$ distillation rounds $r\in\{1,2,\dots,R\}$, each using $C_{r}$ copies of a distillation unit $M_{r}$ operating at code distance $d_{r}$. We assume a negligible cost of state injection and of transferring the output magic states from one level of distillation to the correct positions and code distances to use in the next level. Though each of the mentioned parameters can be optimized, we make multiple a priori decisions regarding them. We assume $R=2$ levels of distillation where $C_{2}=1$, and we assume that $M$ is a 15:1 distillation unit that has been compiled to the same local instruction set that we use in the LSC (see Fig. 9 and Table VI of Ref (Beverland et al., 2022b) for details). We constrain $d_{2}$ to be same code distance as that used in the logical part of the computation (i.e. the full code distance). For a given choice of $d_{1}$, we require that $C_{1}$ be the smallest number that yields $P_{1}^{acc}(\geq 15)\times P_{2}^{acc}\geq P_{th}$, where $P_{1}^{acc}(\geq 15)$ is the probability of at least 15 output states being accepted at the first level, $P_{2}^{acc}$ is the acceptance probability at the second level, and $P_{th}$ is a pre-decided threshold (taken to be 0.985 here). The final run-time of the distillation factory is $\tau(\mathcal{D})\equiv\lceil\tau_{1}+\tau_{2}\rceil$, where $\tau_{r}$ is the number of full-distance slices required for distillation round $r$ [^19] Where $d_{1}\neq d_{2}$, $\tau_{1}=\frac{d_{1}}{d_{2}}\tau_{2}$ may not be a whole number.. The total spatial cost of the distillation factory is $n(\mathcal{D})\equiv\lceil\max_{r}\left[C_{r}n_{r}\right]\rceil=\lceil\max{\left(C_{1}n_{1},n_{2}\right)}\rceil$, where $n_{r}$ is the number of full-distance tiles required for $M$ at distillation round $r$ [^20] Similarly, we calculate $n_{1}$ as $n_{1}=\frac{2d_{1}^{2}-1}{2d_{2}^{2}-1}n_{2}$. The active volume of the distillation factory is approximated to be $V(\mathcal{D})\simeq C_{1}n_{1}\tau_{1}+n_{2}\tau_{2}$. In our resource analysis script, we treat the factory selection process as a function that receives $d_{1}$ and $d_{2}$ as parameters and outputs $n(\mathcal{D})$, $\tau(\mathcal{D})$, $V(\mathcal{D})$, and the distilled state error rate $P_{T}(\mathcal{D})$. This function is used as a subroutine within the resource optimization process to be described.
@@
The object is to calculate $N$, which is the number of magic states that need to be produced in each distillation cycle to keep up with the consumption rate. We note that, since our distillation factories output one state each distillation cycle, $N$ is equal to the number of factories. To cover the requests made during the first cycle, we require $w=\lceil m(1)/N\rceil$ warm-up distillation cycles. Then, we minimize $N$ with respect to the condition $m(k)\leq N(w+k-1)$ because the number of consumed states after $k$ distillation cycles has to be less than or equal to the number produced as of the $(w+k-1)$th cycle at every $k$. Using the expression for $w$, we relax this condition to $m(k)\leq m(1)+N(k-1)$ and solve for N:
(2) $N\geq\frac{m(k)-m(1)}{k-1}$
@@
(3) $N=\left\lceil\max_{k}\frac{m(k)-m(1)}{k-1}\right\rceil.$